# CoLA: Compute-Efficient Pre-Training of LLMs via Low-Rank Activation

### v1

@@ 4.2 CoLA Enables Efficient Checkpointing
Methods Memory Saving Re-Compute
LLAMA-1B w/ vanilla GCP 20.25 GB 1×1$\times$1 ×
CoLA-M-1B 18.94 GB 0.22×0.$2$
Table 5: Quantitative comparison of memory saving and re-computation between vanilla GCP and CoLA-M.
@@
Although delicate optimizations of GCP is beyond our scope, we show in Table 5 and in Fig. 7 the quantitative results and scaling behavior of GCP on LLaMA-1B when applying a heuristic checkpointing strategy. We observe that CoLA-M greatly reduces re-computation cost while achieving similar memory saving as vanilla GCP.
@@ 5 Experiments
CoLA-M 28.82 22.76 16.21 14.59 13.82
Table 7: Validation perplexity of LLaMA-7B pre-trained on C4 dataset. 8-bit Adam/GaLore are collected from Zhao et al. (2024). SLTrain is collected from Han et al. (2024). No results of BF16 Adam reported. We emphasize that CoLA outperforms other methods at 150k steps (14.61 & 14.65) by using only 65k steps.
5.1 Pre-Training LLaMA on C4
We validate our proposed methods by extensively pre-training LLaMA-like LLMs from 60M to 7B scales. Experiments were performed on NVIDIA A100/H100 GPUs. We closely follow the experiment settings of Zhao et al. (2024); Han et al. (2024), and directly compare CoLA with their reported results. We use the C4 dataset Raffel et al. (2020), which is a colossal, cleaned version of Common Crawl’s web crawl corpus. C4 dataset has been widely used for pre-training LLMs. Trainings were done without data repetition on a sufficiently large amount of tokens. We compare CoLA with baselines including full-rank pre-training, ReLoRA Hu et al. (2021), GaLore Zhao et al. (2024), and SLTrain Han et al. (2024), with a focus on methods that explore model efficiency.
We implement CoLA and its memory efficient variant CoLA-M by parameterizing all linear layers into the proposed linear-nonlinear-linear composition [i.e. Eq. (3)], and keep all other parameters and operations unchanged. We use AdamW optimizer and cosine annealing learning rate scheduler Loshchilov and Hutter (2016) with warm-up. We remark that CoLA is NO more sensitive to optimizer-related hyper-parameters. We refer more details to Appendix D.
Table 6 compares our methods and other efficient pre-training techniques in terms of validation perplexity, parameter size, and estimated memory usage of model, gradients and optimizer states. CoLA has the smallest model size, thereby consumes the least memory, and demonstrates on-par validation perplexity compared to the full-rank baselines. Table 7 compares the validation perplexity on the 7B model. We highlight the fact that CoLA outperforms 8-bit Adam/GaLore at their 150k steps (14.61 & 14.65) using only 65k steps. Meanwhile the total memory cost of CoLA-M is less than half of the other methods. We also emphasize our comparison with GaLore and SLTrain. Our proposed method has an overall better performance, whilst further reduces model size on top of the reduction of SLTrain. More importantly, we follow our discussion in Section 3.3 and remark that CoLA has uniformly fewer computations than GaLore and SLTrain when applying the same rank.
5.2 Scaling Behavior
@@
We briefly discuss how CoLA might scale differently compared to full-rank training. A comprehensive investigation of this topic is beyond the scope of this work due to the huge computing cost.
Table 8 shows a few experiments on how CoLA might be improved when computing is scaled up. The default rank choice reduces the compute cost to about a half of full-rank training, without significantly hurting the model performance. Meanwhile, if we relax the computing restriction and increase the rank to be greater than one quarter of $d$, then CoLA outperforms full-rank training in all three scales, while still being able to reduce the computing cost. One might argue that full-rank training can also be scaled down to a similar compute of CoLA and might perform similarly. We implement such baselines in Table 8 and refer this setup to “Control". We typically reduce the number of layers or the model width of full-rank models to scale down their computing cost. We find empirically that they will tend to reduce performance quickly and dramatically underperform CoLA.
5.3 System Measurements
@@
Mem (GB) Throughput FLOPs Mem (GB) Throughput FLOPs
@@
We further investigate the efficiency of CoLA from a more practical perspective. It is often observed that a theoretically less expensive method can have worse system-level performance due to poorly designed hardware implementation or lack of system-aware optimizations. We show that this is NOT the case for CoLA, by illustrating that its out-of-the-box system performance already significantly outperforms the full-rank training and other efficient training methods. We focus on the actual memory usage and the pre-training throughput.
Fig. 8 shows the measured throughput for pre-training the LLaMA/CoLA 1B model. The sequence batch size is set as 16, which fully utilizes the A100 GPU. Among all these methods, CoLA and CoLA-M are the only two that show higher GPU throughput than the full-rank baseline, while all other methods downgrade throughput due to their computing overhead. In particular, CoLA-M, the memory-efficient CoLA that significantly reduces overall GPU memory, still shows higher training throughput despite the re-computation overhead. Meanwhile, vanilla GCP, which uses a similar idea of trading compute for memory, reduces throughput from the full-rank baseline by 26%. We show further the details of these measurements in Table 9, and also compare their estimated FLOPs. At both 1B and 7B scale, CoLA-M manages to almost halve the computing cost and reduce the memory usage by two thirds, achieving a great balance between computing and memory efficiency. We refer more details of our system profiling to Appendix. F.
5.4 Inference Performance
1B (BZ=32) 7B (BZ=32)
Mem (GB) Throughput Mem (GB) Throughput
Full-rank 5.74 21,109 18.15 11,086
SLTrain 4.18 20,096 12.70 9,968
CoLA 3.84 34,697 10.87 16,012
We highlight the fact that CoLA not only reduces pre-training resources but also speeds up inference and reduces its memory cost. Table 10 shows that CoLA improves inference throughput by up to $\bf 1.64\boldsymbol{\times}$ while reducing memory cost by up to $\bf 1.67\boldsymbol{\times}$.
@@ 6 Conclusions
We propose CoLA, and its memory efficient variant CoLA-M, to achieve collectively parameter, compute and memory efficiency at both pre-training and inference time for large foundation models. CoLA effectively reduces 2×2$\times$2 × model size while preserving full-rank level performance. More importantly, we show the reduction does not come with additional compute. Instead, CoLA halves compute and almost doubles training throughput from its full-rank counterpart. When memory is of higher concerns, CoLA-M trades only minimum compute for state-of-the-art memory reduction during pre-training, meanwhile still reducing compute and improving throughput. We hope our work will inspire the community to further investigate the architecture efficiency that has been overlooked and under-discovered for large foundation models.
@@ 7 Limitations
This work limits the study of our proposed formulation under LLaMA-like architectures. The adaptation of CoLA to other architectures is conceptually trivial, but their performance is to be evaluated via real experiments. In this work, we only pre-train each model to be roughly compute-optimal (for original LLaMA models, not CoLA), while industry-produced LLMs (that are of similar scales) are often over-trained. It is worth investigating the performance of CoLA when significantly over-trained. We leave this computing-expensive research for future work.
@@ Appendix D Hyper-Parameters
Appendix E Ablation Study

### v2

@@ 4.2 CoLA Enables Efficient Checkpointing
Although delicate optimizations of GCP is beyond our scope, we show in Fig. 7 the quantitative results and scaling behavior of GCP on LLaMA-1B when applying a heuristic checkpointing strategy. We observe that CoLA-M greatly reduces re-computation cost by $\bf 4.6\boldsymbol{\times}$ while achieving similar memory saving (18.94GB) as vanilla GCP (20.25GB).
@@ 5 Experiments
CoLA-M 26.82 22.76 16.21 13.82 13.09 12.73
Table 6: Validation perplexity of LLaMA-7B pre-trained on C4 dataset. 8-bit Adam/GaLore are collected from Zhao et al. (2024). SLTrain is collected from Han et al. (2024). No results of BF16 Adam reported.
@@
Pre-Training Loss QQP SST-2 MRPC COLA QNLI MNLI RTE STS-B GLUE Avg
[…]
CoLA 1.257 91.2 92.3 90.6 54.1 91.7 84.3 74.2 89.7 83.5
[…]
5.1 Pre-Training within Compute-Optimal
We validate our proposed methods by extensively pre-training LLaMA-like LLMs from 60M to 7B scales following the exact experimental setup in Zhao et al. (2024); Han et al. (2024). Trainings were done using C4 dataset Raffel et al. (2020) without data repetition on roughly compute-optimal [^2] Compute optimal regime refers to the token-to-parameter (T2P) ratio being ~20 Hoffmann et al. (2022). amounts of tokens. We compare CoLA with baselines including full-rank pre-training, ReLoRA Hu et al. (2021), GaLore Zhao et al. (2024), and SLTrain Han et al. (2024), with a focus on methods that explore model efficiency.
We implement CoLA and CoLA-M by parameterizing all MLP layers and all projection layers in attention with auto-encoders [i.e. Eq. (3)], and keep all other parameters and operations unchanged. We use AdamW optimizer and cosine annealing learning rate scheduler Loshchilov and Hutter (2016) with warm-up. We refer detailed configurations to Appendix D.
Table 5 compares our methods and other efficient pre-training techniques in terms of validation perplexity, parameter size, and estimated memory usage of model, gradients and optimizer states. CoLA has the smallest model size, thereby consumes the least memory, and performs on-par with full-rank baselines. CoLA uniformly surpasses other efficient training baselines in both efficiency and performance. Table 6 compares the validation perplexity on the 7B model for 150k steps [^3] Due to resources constraints, 7B models are trained below compute optimal budget Zhao et al. (2024); Han et al. (2024).. CoLA(-M) significantly outperforms 8-bit Adam/GaLore by 12.73 vs ~14.6, while saving two-third memory.
Scaling Behavior: Table 7 shows how CoLA might be improved when compute is scaled up. The default rank choices reduce half the computing cost, without harming the model performance. Meanwhile, if we relax the computing restriction and moderately increase the rank, then CoLA outperforms full-rank training in all three scales, while still being fairly smaller and reducing the computing cost. One might argue that full-rank training can also be scaled down to a similar computing cost of CoLA and might perform similarly. We implement such baselines in Table 7 and refer this setup to “Control". We typically reduce the number of layers or the model width of full-rank models to scale down their computing cost. We find empirically that they reduce performance significantly and dramatically underperform CoLA.
5.2 Pre-Training beyond Compute-Optimal
According to Chinchilla scaling law Hoffmann et al. (2022), compute-optimal training is at the efficient frontier when given a fixed computing budget or a target model size. However, leading industrial groups with massive computing resources tend to extensively overtrain smaller models for efficient deployment, such as LLaMA-3 Grattafiori et al. (2024) 1-3B models being trained up to 9 Trillion tokens. To evaluate CoLA’s effectiveness beyond the compute-optimal regime, we further experiment the following two over-training settings.
LLaMA-350M with 51B Tokens: We prolong the training duration by 8$\times$ of the compute-optimal budget for both CoLA [^4] We choose CoLA at 0.7×0.7$\times$0.7 × compute of full-rank baseline, as its superior performance observed in Table 7. and full-rank LLaMA at 350M scale. This results in 51B total training tokens. CoLA continues outperforming full-rank baseline on validation perplexity of 13.96 vs 14.47, consistent with results at compute-optimal observed from Table 7.
[…]
These results further demonstrate CoLA’s effectiveness across both encoder/decoder architectures, both compute-optimal/over-train settings, and different activations (GeLU, Swish and SwiGLU).
5.3 Training/Inference System Performance
@@
Mem (GB) Token/s FLOPs Mem (GB) Token/s FLOPs
@@
Superior Training Efficiency.
We further validate CoLA’s efficiency from a practical perspective: CoLA delivers superior out-of-the-box system performance compared to full-rank and other efficient training methods. Fig. 8 compares pre-training throughput for the 1B-scale LLaMA model (batch size 16, fully utilizing A100 GPUs). Among evaluated methods, only CoLA and CoLA-M surpass the full-rank baseline throughput. Notably, CoLA-M maintains higher throughput despite recomputation overhead, significantly outperforming vanilla GCP. Table 9 provides detailed measurements, showing CoLA-M cuts computing cost nearly by half and reduces memory usage by two-thirds, achieving great balance between memory and compute efficiency. Profiling details are available in Appendix F.
Superior Inference Efficiency.
Not just for training, CoLA also speeds up inference and reduces memory cost. Table 11 (Appendix E.2) shows that CoLA off-the-shelf improves inference throughput by up to $\bf 1.64\boldsymbol{\times}$ while reducing memory cost by up to $\bf 1.67\boldsymbol{\times}$.
@@ 6 Conclusions
We have proposed CoLA, and its memory efficient variant CoLA-M, to achieve collectively parameter, computing and memory efficiency at both training and inference time for large foundation models. CoLA effectively reduces $\bf 2\boldsymbol{\times}$ model size and computing cost while preserving full-rank level performance. CoLA-M trades minimum overhead for state-of-the-art memory reduction, while still improving training throughput over full-rank baselines. Crucially, CoLA is promising to save substantial GPU resources in LLM industry. This work has been focused on dense architectures. In the future, it is worth extending CoLA to the mixture-of-expert (MoE) architecture.
@@ 7 Limitations
[…]
@@ D.1 LLaMA Pre-Training
[…]
Appendix E Additional Results
E.1 Ablation Study
@@
E.2 Inference Efficiency
1B (BZ=32) 7B (BZ=32)
Mem (GB) Token/s Mem (GB) Token/s
Full-rank 5.74 21,109 18.15 11,086
SLTrain 4.18 20,096 12.70 9,968
CoLA 3.84 34,697 10.87 16,012
We show CoLA’s system performance at inference stage in Table 11. CoLA reduces memory usage and improves inference throughput compared to full-rank baselines.